OpenStudy (anonymous):

An Olympic bobsled starts from rest and slides down a frictionless 9.00 slope. a) what is the final velocity after 18 seconds?

OpenStudy (anonymous):

you mean that theta=9?

OpenStudy (anonymous):

\[mgsin \theta=ma \rightarrow a=gsin \theta \rightarrow V=(gsin \theta)t\]

OpenStudy (anonymous):

yeah \[\phi= 9 degrees\]

OpenStudy (anonymous):

so put theta in above friend

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

so \[V=(g \sin \theta) t\] will be V=(9.8sin(9))18= 27.59

OpenStudy (anonymous):

m/s

OpenStudy (anonymous):

yes i got it correct..thank you

OpenStudy (anonymous):

your wellcome